OpenStudy (anonymous):

Given that f(x)= 4x-5 and g(x)=x^3+1, find (fog)(-1)

OpenStudy (anonymous):

You're expression says F of G of X so first you must solve your G(x) with the given substitute for X and then use the result to solve the F(x)

OpenStudy (anonymous):

\[F(x) = 4x-5\] \[G(x) = x^3+1\] And initially you are given F(g(-1)) so use the -1 to solve for G(x) then from the result of g(x) use it to solve for f(x). Hence "F of G of X or simply F(g(x))

OpenStudy (anonymous):

I don't understand pls

OpenStudy (anonymous):

g(x) is an expression in terms of X and it is telling you that X will be -1 so plug that into your X for G(x)

OpenStudy (anonymous):

Replace all the "x" values in your expression G(x) with (-1), very trivial. \[G(-1) = (-1)^3 +1\]

OpenStudy (anonymous):

f(x)=4(-1)-5=-9?

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

Solve the G(x) first since it is on the outside, No that is incorrect look above.

OpenStudy (anonymous):

F of G of X means first G(x) then using the result/answer that you get plug it into F(x)

OpenStudy (anonymous):

So solve the G(x) first by using -1 as x then you will have another number and use that number in the f(x) and you will find your answer.

OpenStudy (anonymous):

(-1)^3+1=0

OpenStudy (anonymous):

Correct! Now use that value in place for the "x" in F(x)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(x)=4x-5= 4(0)-5= -5

OpenStudy (anonymous):

Correct! Great Job

OpenStudy (anonymous):

Now that is your f(g(x))

OpenStudy (anonymous):

Is that all @Johnbc

OpenStudy (anonymous):

Thanks for your help @Johnbc

OpenStudy (anonymous):

Yes you have found your answer, it has been a pleasure.